OpenStudy (anonymous):

what are the solutions of x^2-3x=-2x-7?

OpenStudy (callisto):

x^2-3x=-2x-7 x^2 -x +7 =0 Time to use quadratic formula~

sam (.sam.):

x^(2)-3x=-2x-7 Since -2x contains the variable to solve for, move it to the left-hand side of the equation by adding 2x to both sides. x^(2)-3x+2x=-7 Since -3x and 2x are like terms, subtract 2x from -3x to get -x. x^(2)-x=-7 To set the left-hand side of the equation equal to 0, move all the expressions to the left-hand side. x^(2)-x+7=0 Use the quadratic formula to find the solutions. In this case, the values are a=1, b=-1, and c=7.

OpenStudy (callisto):

\[x= \frac{-(-1)\pm \sqrt{(-1)^2 - 4(1)(7)}}{2(1)}\]\[x= \frac{1\pm \sqrt{-27}}{2} =...\]

OpenStudy (anonymous):

thank you